if A=2B then b(b+c)=
 a) a2  b)a  c)2a   d)a/2 (from properties of triangles)
 

Grade:12

1 Answers

moumi roy
6 years ago
ANSWER IS (A)
THIS CAN BE DONE BY HIT AND TRIAL
TAKING B=45  A=90.HENCE AN ISOSCALE TRIANGLE where b=c
if b=c=1,a=(2^(1/2))
putting in b(b+c)=2=a^2.check it ,it is true for all.